1. The greatest of four consecutive even integers is 14 less than twice the smalles integer. What are the integers?
2. The degree of measurement of angles of a triangle are consecutive even integers. What is the degree measure of each of the three angles?
3. The sum of two numbers is 18. Three times the greater number exceeds four times the smaller number by 5. What are the numbers?

Answer by longjonsilver(2297)

1. Let smallest number be x. So next one is x+2, then x+4 and largest is x+6

x+6 = 2(x) - 14
x+6 = 2x - 14
6 = x - 14
20 = x

so numbers are 20, 22, 24, 26

2. Let smallest angle = a
next angle is a+1
largest is a+2

All 3 add to 180 degrees, so (a+a+1+a+2) = 180
3a+3 = 180
3a = 177
a = 177/3
a = 59

so 3 angles are 59, 60 and 61 degrees

3. Let 2 numbers be x and y, with x being the larger.

x+y = 18
3x - 4y = 5

You solve this simultaneous equation.

jon.